OWG Meeting 10-Sep-2015

From GlueXWiki
Jump to: navigation, search

Location and Time

Room: CC F326-327

Time: 1:30pm-2:30pm


You can connect using BlueJeans Video conferencing (ID: 120390084). (Click "Expand" to the right for details -->):

(if problems, call phone in conference room: 757-269-6460)

  1. To join via Polycom room system go to the IP Address: (bjn.vc) and enter the meeting ID: 120390084.
  2. To join via a Web Browser, go to the page https://bluejeans.com/120390084.
  3. To join via phone, use one of the following numbers and the Conference ID: 120390084
    • US or Canada: +1 408 740 7256 or
    • US or Canada: +1 888 240 2560
  4. More information on connecting to bluejeans is available.

Previous Meeting


  1. Announcements
  2. GlueX Collaboration Meeting October 8-10
  3. Upgrade to RHEL7
    • Thur. Sept. 17
  4. ROL modifications
    • SYNC events
    • TS info for physics events (L1 latch word, unix timestamp)
  5. Readout Config. parameters in CCDB (RCDB?)
  6. L3 development
  7. DAQ system development
  8. AOT


Attendees: David L. (chair), Simon T., Sergey F., Dave A., Hovanes E., Curtis M., Sean D., Vardan G., Alex S., Beni Z., Carl T.


  • Solenoid work plan anticipates it will not be cooled down until early Jan.
  • There is still a possibility of running shifts with zero field in Dec. (not in Nov.). Possible uses for beam with no solenoid:
    • L1 Trigger Commissioning
    • Diamond commissioning (unclear if they will be available then)
    • Tagger, PS, TPOL, TAC commissioning
    • DAQ commissioning
  • There will need to be a collaboration wide discussion on this

Collaboration Meeting

  • We will fill in same 4 talks as last meeting with these changes
    1. Alex's L1 trigger talk will include some details of the new fADC250 firmware specs.
    2. Add a talk on the fADC125 (Naomi was unanimously assigned to give this)


  • There is a preliminary firmware version for the F1 TDC that that addresses the issue of failing at high rates
    • Alex has done some testing using a single crate, but has seen rates put to 60kHz
    • Ed is still tweaking the firmware but the timescale for a final version is unknown
    • There was some discussion on the time scale for doing a multi-crate test using the preliminary firmware we have. Alex had some slight reservations about deploying it widely since it requires a different library. This only complicates the process of updating the firmware in many modules/crates for the test.
    • We will discuss it some more offline to see how feasible it is for Sergey to do multi-crate testing with the preliminary firmware version prior to the Oct. collaboration meeting.
  • Other debugging efforts are ongoing with no significant milestones to support.
    • Ironing out minor wrinkles in CODA 3.04
    • Testing and debugging of intermediate version of fADC125 for high rates
  • Sergey continues work on a pedestal monitoring system that runs on the ROCs


  • After some discussion, we decided to defer the upgrade to RHEL7 until after the next run. We will likely revisit it next summer. Some of the highlights of the discussion:
    • The version of the Lustre file system currently used for the volatile and cache disks on the CUE does not have a driver available for RHEL7.
    • The JLab CC has plans to migrate the work disks over to this format as well
    • Recent fixes [1][2][3]) have removed some motivation for the upgrade.
    • While gluon46 has not been spontaneously rebooting since being upgraded to RHEL7, it is believed that this would have also been the case if RHEL6 had been reinstalled.
  • We decided to keep gluon01 and gluon46 running RHEL7 but leave all others as they are.
    • We will ask that the Hall-D channel be moved from RHEL 7.0 to 7.1 and that gluon46 be rebuilt using the channel. Currently, gluon01 is on the Hall-D channel and running 7.0. gluon46 is not on the channel and is running 7.1. This will bring them into alignment and on 7.1. They will serve as testing platforms until we decide to upgrade.

ROL Modifications

  • SYNC events
    • Alex has done some testing/development using a couple of crates (including the TS crate)
    • New options related to SYNC events have been added to the DAQ configuration files
    • Next steps will involve adding asynchronous pedestal information to data stream
    • DAC will not be automatically updated in initial version but that functionality can be added later
  • TS info block may be uncommented already, but no one recalls doing it. Alex will look at it again.
  • Alex has made some modifications to the TS library. He agreed to send these to Bryan M. so that they may be incorporated into the official version.


  • No progress here
  • Dmitry's status regarding visiting JLab is unchanged as far as anyone knows
  • Sergey communicated with Dmitry last week and said he has plans to work on the RCDB web interface issue remotely

L3 Development

  • Nothing to report